OpenStudy (totallmarcus):

What is the solution to the following system of equations? y = −x2 − 5x − 4 y = −x2 + 9x − 18

OpenStudy (triciaal):

set the y's equal and solve the equation to get the x-value. use the x value in either expression for y to get the corresponding y-value.

OpenStudy (bunny12):

Yes in other words, substitute the y

OpenStudy (blowery):

-x^2 - 5x - 4 = -x^2 + 9x - 18 The negative x squared will cancel out leaving -5x - 4 = 9x - 18 Next, move the -5x to the other side of the equal mark, -4 = 14x - 18 Next, move the -18 to the other side of the equal mark leaving 14 = 14x Divide through by 14 leaving 1 = x Now, substitute the 1 for x in the equation y = -x^2 - 5x - 4 Y = -1^2 - 5(1) - 4 y = 1 - 5 - 4 y = -8

OpenStudy (triciaal):

be careful with the simple stuff -1-5-4 = -9